# Large Language Model-Enabled Multi-Agent Manufacturing Systems

### v1

@@ IV-B Case Study Simulation and Results Analysis
Following the setup of our case study, simulations were conducted to evaluate the system performance, as shown in one example, depicted in Figure 5. In this example, the user initiates the workflow by instructing the ‘Product-1’ agent and all other RAs, using  Algorithm 1 from Section III-A. Following the agent initialization, the system operates according to the steps detailed in the operational workflow, as depicted in Section III-C. First, the ‘Product-1’ agent communicates the need for contouring to the ’Milling1’ agent, providing only the contouring section of G-code specifications. After the ‘Milling1’ agent completes the task, it informs the ‘Product-1’ agent, indicating it is ready for the next operation. The process continues with the ‘Product-1’ agent assigning the drilling operation to the ‘Drilling1’ agent, which updates the ‘Product-1’ agent after completion. When all operations are completed, the ‘Product-1’ agent updates the user, indicating that all manufacturing processes for ‘Product-1’ are completed and are now ready to move to the exit buffer. All agent communications are facilitated by the chat interaction protocol defined in Algorithm 2, as outlined in Section III-B.
This section details the outcomes from the simulations, highlighting the communications among agents, their understanding and execution of tasks, the system’s accuracy, and the insights gained from these case studies.
@@ IV-B1 Communication and Comprehension
Our case study focused on evaluating the agents’ ability to communicate and comprehend specific tasks within the manufacturing process. Utilizing natural language processing, agents interacted in both the 2-step and 4-step processes, showcasing their capability for nuanced communication. In the 2-step process, the communication between the ’Product-1’ agent and the ‘Milling1’ and ‘Drilling1’ agents showed the system’s efficiency in handling straightforward task sequences and distributing G-code specifications in human-like communication. The 4-step process further explored the agents’ capability in managing and understanding complex sequences, requiring accurate G-code allocation where ‘Milling1’ was tasked with both contouring and counterboring. In this scenario, the ‘Product-1’ agent coordinated the sequence of contouring, drilling, counterboring, and threading, showing the agents’ ability to interpret and implement longer instructions.
@@ IV-B2 Accuracy and Performance
The data presented in Table I reveals that GPT-4 achieved success rates of 100% for 2-step and 86% for 4-step manufacturing processes across 50 trials for each. This shows a decline in performance as task complexity increases, with the 4-step process experiencing a 14% overall error rate. The detailed error breakdown in  Table II reveals that within the overall 14% error rate for the 4-step process, incorrect function calls and process inexecution have 43% of the errors, while inaccurate G-code allocation accounts for 14% of the total errors. It is observed that the increase in prompt length correlates with a rise in errors, highlighting the need for refined model training and improved prompt engineering. Addressing these issues aims to enhance task execution capabilities and effective integration of LLMs into MASs for manufacturing, improving system adaptability and operational efficiency even in unexpected scenarios.
@@ IV-B3 Insights from Case Studies
The case study provided insights into the capabilities of LLM-enabled MASs in manufacturing. The current framework incorporates LLM to transition towards a system capable of understanding the manufacturing tasks (Cn1), interpreting and adapting to new product specifications (Cn2), and agent communication via natural language instructions (Cn3). This shift enhances system flexibility and expands its ability to manage various tasks without relying on predetermined models or rules. However, accuracy issues with GPT-4 highlight the need for model improvement and adaptation to manufacturing needs. Difficulties with incorrect argument passing during function calls point to the necessity for better prompt engineering, error management, and validation within the system. These results guide future research toward enhancing LLM integration in manufacturing, developing advanced agent communication protocols, and increasing system robustness against errors.
@@ V-A Performance
A significant challenge within the use of LLMs in manufacturing processes is achieving the level of performance required for high-quality production. Manufacturing processes demand precision to ensure product consistency, safety, and reliability. However, LLMs, while effective in processing natural language, occasionally shows limitations in delivering the precision necessary for these applications, as indicated in Section IV-B2. This issue raises concerns about the feasibility of relying on LLMs for tasks where errors can lead to other consequences, such as machine breakdown and low product quality.
@@ V-B Scalability
Scalability is also a challenge when integrating LLMs into manufacturing systems. As the volume of data and the complexity of tasks grow, there is a risk that LLMs may not maintain their initial levels of accuracy and efficiency. This limitation is particularly noticeable in situations where extensive historical data and detailed prompts are essential for task execution, as exemplified in the 4-step process in Section IV. The potential decrease in accuracy with larger datasets poses a challenge to scaling LLM-enabled multi-agent control across more complex manufacturing operations.
@@ V-C Reliability
The integration of LLMs into manufacturing settings highlights the importance of reliability. LLMs generating hallucinated responses or making false decisions raises significant safety concerns, as shown in Section IV. The LLM in the case study allocated inaccurate G-code or executed the incorrect function calls. Moreover, the reasoning provided by LLMs may require validation in practical manufacturing scenarios even if it seems logical. Therefore, it is essential to implement robust verification mechanisms, as well as incorporate real-time observability and monitoring capabilities. This approach ensures LLMs’ outputs and reasoning are consistent and allow immediate adjustments with manufacturing standards and safety protocols. For instance, the study by Joublin et al. [35] presents a system architecture for robotics that utilizes LLMs for planning tasks, focusing on a strategy for replanning in response to errors via feedback. This method enables robots to adjust to failures and changes, enhancing reliability. Thus, feedback loops in LLM systems ensure safety and reliability in complex manufacturing environments.
@@ VI Conclusion and Future Work
In our research, we explored the integration of LLM within MAS to enhance manufacturing processes. Our research demonstrates the effect of LLM in manufacturing, where they serve to dynamically match tasks with optimal process parameters, improving task allocation efficiency and operational adaptability (Cn1). Moreover, LLM enables agents to adapt to evolving product specifications (e.g. G-code), meeting the manufacturing requirements (Cn2). Additionally, through advanced natural language processing, LLM enables agents to effectively understand and execute complex, context-specific instructions from humans and other agents enhancing the agent communication and collaboration within the manufacturing process (Cn3). However, challenges have been identified, such as task execution performance, scalability, and reliability, highlighting the need for a verification process and monitoring capabilities when making decisions for manufacturing operations. Furthermore, the diversity in user instructions based on different skill levels and manufacturing requirements affects task interpretation and execution, which requires a strategy to dynamically adapt to these changes when integrating LLM to ensure reliability.
Future work will focus on developing communication protocols incorporating LLM for agents that can accommodate a wide variety of manufacturing instructions and requirements, influenced by varying proficiency levels of human operators. This approach seeks to showcase the capability of agents to learn and adapt from historical data, enhancing their responsiveness to rapid changes in modern manufacturing industries. Additionally, efforts will be directed towards enhancing the accuracy and scalability of LLM. While this paper initially focused on widely used manufacturing processes (e.g. drilling, milling), future investigations will extend to include less commonly used manufacturing processes with larger datasets to evaluate the effectiveness of our framework. The goal is to assess if integrating additional processes results in less effective outcomes or if they can also be improved by applying these methodologies, ensuring efficient processing of larger datasets and diverse manufacturing processes without impacting performance.

### v2

@@ IV-B Case Study Simulation and Results Analysis
Following the case study setup, simulations were conducted to evaluate the system, as shown in Figure 5. In this example, the user initiates the workflow by instructing the ‘Product-1’ agent and all RAs, using  Algorithm 1 from Section III-A. After initialization, the system follows the operational workflow in Section III-C. The ‘Product-1’ agent first communicates the need for contouring to the ’Milling1’ agent, providing only the contouring G-code section. After the ‘Milling1’ agent completes the task, it informs the ‘Product-1’ agent, indicating it is ready for the next operation. The process continues with the ‘Product-1’ agent assigning the drilling operation to the ‘Drilling1’ agent, which updates the ‘Product-1’ agent after completion. When all operations are completed, the ‘Product-1’ agent updates the user, indicating it is ready to move to the exit buffer. All agent communications are facilitated by the chat interaction protocol defined in Algorithm 2, as outlined in Section III-B.
@@ IV-B1 Communication and Comprehension
Our case study evaluated the agents’ ability to communicate and comprehend specific tasks. Using natural language processing, agents interacted in both 2-step and 4-step processes. In the 2-step process, the communication between the ’Product-1’ agent and the ‘Milling1’ and ‘Drilling1’ agents showed the system’s efficiency in handling straightforward task sequences and distributing G-code specifications. The 4-step process tested the agents’ ability to manage complex sequences, with ‘Milling1’ handling both contouring and counterboring. In this scenario, the ‘Product-1’ agent coordinated contouring, drilling, counterboring, and threading, showing the agents’ ability to interpret complex instructions.
@@ IV-B2 Accuracy and Performance
The data in Table I shows GPT-4 achieved 100% success in 2-step and 86% in 4-step processes across 50 trials each, with performance declining as task complexity increased. The 4-step process had a 14% overall error rate. The detailed error breakdown in Table II shows that in the 4-step process, 43% of errors were due to incorrect function calls and process failures, and 14% were due to inaccurate G-code allocation. The increase in prompt length correlates with a rise in errors, highlighting the need for model training and improved prompt engineering. Addressing these issues will enhance task execution capabilities in MASs for manufacturing, improving system adaptability and efficiency even in unexpected scenarios.
@@ IV-B3 Insights from Case Studies
The case study provided insights into the capabilities of LLM-enabled MASs in manufacturing. The framework uses LLM to understand manufacturing tasks (Cn1), adapt to new product specifications (Cn2), and facilitate agent communication via natural language (Cn3). This enhances system flexibility and task management without relying on predetermined rules. However, accuracy issues with GPT-4 highlight the need for model improvement to manufacturing needs. Difficulties with incorrect function calls point to the necessity for better error management and validation. These results guide future research toward enhancing LLM integration, developing advanced agent communication protocols, and increasing system robustness.
@@ V-A Performance
A significant challenge in manufacturing using LLMs is achieving the performance required for high-quality production. While effective in processing natural language, LLMs sometimes lack precision, as indicated in Section IV-B2. This issue raises concerns about the feasibility of relying on LLMs for tasks where errors can cause machine breakdown and low product quality. The proposed framework can enhance performance by applying fine-tuned LLMs on domain-specific data [32].
@@ V-B Scalability
As data volume and task complexity grow, LLMs may struggle to maintain accuracy. This limitation is particularly noticeable in situations where extensive historical data and detailed prompts are essential, as exemplified in the 4-step process in Section IV. The potential decrease in accuracy with larger datasets poses a challenge to scaling LLM-enabled multi-agent control across more complex manufacturing operations. A knowledge base for multi-agent manufacturing systems [33] should be integrated for practical industrial applications to provide LLMs with accurate runtime data and to enhance scalability.
@@ V-C Reliability
Reliability is key for integrating this method into real manufacturing scenarios. LLMs generating hallucinated responses or making false decisions raises safety concerns, as shown in Section IV. The LLM allocated inaccurate G-code or executed the incorrect function calls. Moreover, the reasoning provided by LLMs may require validation in practical manufacturing scenarios even if it seems logical. Therefore, it is essential to implement robust verification mechanisms, as well as incorporate real-time observability and monitoring capabilities. This approach ensures LLMs’ outputs and reasoning are consistent and allow immediate adjustments with manufacturing standards and safety protocols. Joublin et al. [34] demonstrate a system for robotics that uses LLMs for task planning and replanning in response to errors, highlighting the importance of feedback loops for safety and reliability in complex manufacturing environments.
@@ VI Conclusion and Future Work
In our research, we explored integrating LLM within MAS to enhance manufacturing processes. Our research demonstrates that LLM improves task allocation efficiency and operational adaptability (Cn1), enables agents to adapt to evolving product specifications (e.g., G-code) (Cn2), and enhances agent communication and collaboration through advanced natural language processing (Cn3). However, challenges like task execution performance, scalability, and reliability highlight the need for verification and monitoring. Additionally, the diversity in user instructions requires a strategy to ensure reliable task interpretation and execution.
Future work will focus on developing LLM-based communication protocols for agents to handle diverse manufacturing instructions and operator proficiency levels. This approach seeks to showcase the agents’ capability to learn and adapt from historical data, enhancing their responsiveness to rapid industry changes. Additionally, efforts will be directed towards enhancing the accuracy and scalability of LLM. While this paper focused on common manufacturing processes (e.g. drilling, milling), future research will include less common processes with larger datasets to evaluate the framework’s effectiveness. The goal is to ensure the efficient processing of diverse manufacturing tasks without impacting performance.